OpenStudy (anonymous):

PLEASE HELP Find the solution to the system of equations by substitution. p = q + 2 2p + 3q = -26

OpenStudy (chaise):

p = q + 2 Eqn(1) 2p + 3q = -26 Eqn(2) With these sets of equations, we can substitute the value of p in equation 1, in to equation 2, with one equation and one unknown, we can begin to solve for q. 2(q+2)+3q=-26 Expanding the brackets gives 2q+4+3q=-26 Collecting like terms: 5q+4=-26 Subtracting 4 from both sides 5q=-30 Divide 5 from both sides q=-6

OpenStudy (chaise):

This value of q can be used to find p in equation 1.

OpenStudy (anonymous):

ty